Solve for x
x with a 2 as the exponent - 6x + 9 = 0

Answer by checkley71(8403)

X^2-6X+9=0
(X-3)(X-3)=0
X-3=0
X=3 AMNSWER.
X-3=0
X=3 ANSWER.